OpenStudy (anonymous):

If f(x)=5x^2-x+2 Then, f(x+h)-f(x)/h =?

hartnn (hartnn):

what did u get f(x+h) as ?

OpenStudy (anonymous):

I'm stuck at 5(x+h)^2-(x+h)+2

OpenStudy (anonymous):

I dont know what to do with the 5(x+h)^2

hartnn (hartnn):

\( (a+b)^2 = a^2+2ab+b^2 \)

OpenStudy (anonymous):

and then multiply them all by 5??

hartnn (hartnn):

yes!

OpenStudy (anonymous):

the answer i got was wrong

OpenStudy (anonymous):

i got 5x^2+10xh+5h^2-x-h+2

OpenStudy (anonymous):

is that part right?

hartnn (hartnn):

thats correct :) carry on

OpenStudy (anonymous):

5x^2+10x+5h-x+2

hartnn (hartnn):

nopes, its not -f(h) its -f(x)

hartnn (hartnn):

\(f(x+h) = 5 (x+h)^2 = 5 x^2 + 10xh +5h^2 -x-h+2\) \( f(x) = 5x^2-x+2 \) subtract them!

OpenStudy (anonymous):

okay i got 10xh+5h^2-h

OpenStudy (anonymous):

oh wait i got the answer. i put in a sign wrong

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

welcome ^_^